# Analyzing the Generalization and Reliability of Steering Vectors

### v1

@@ Appendix B Limitations
Appendix C Methodological Details
C.1 Contrastive Prompts
@@ C.2 Prompt Template
C.3 Out-of-distribution Prompt Injections
@@
C.4 Aggregation Methods
Besides MD, other aggregation methods have been proposed in the literature, such as principal components analysis (PCA) and logistic regression (LG) [Zou et al., 2023]. Here, we focus on mean-difference as it naturally results in steering vectors which have an appropriate magnitude relative to model activations , such that a steering multiplier of 1 has the same semantic meaning across different layers and different concepts. return steering vectors with unit norm, which could be different from the natural scale of the model. Thus, additional work would be required to tune appropriate steering magnitudes. We stick with MD, as Rimsky et al. [2023] show it is a strong baseline, and PCA has been found to give a very similar direction to MD [Tigges et al., 2023].
@@ C.4 Aggregation Methods
C.5 MCQ vs Open-Ended Generation
For evaluating the effectiveness of steering vectors, the setting of ultimate interest is the open-ended generation setting. However, it is difficult to obtain an objective metric of whether steering vectors are effective in this setting. Previous work [Rimsky et al., 2023, Zou et al., 2023] finds that multiple-choice propensity generally correlates with open-ended propensity. Hence, our analysis primarily focuses on the multiple-choice setting, with examples being prompt-engineered to select one of the multiple-choice options.
C.6 Logit-Difference Propensity
Following standard practice in the mechanistic interpretability literature, we use the difference in logits between a correct and incorrect answer as the metric of propensity. Our use here is justified by two points: (i) Firstly, the correct and wrong answers are unambiguous. We find that, when the prompts are formatted in multiple-choice format, the two highest logits consistently correspond to the option tokens A and B, indicating that it is valid to consider only these two logits. (ii) Secondly, previous work [Rimsky et al., 2023] finds that the logit-difference usually corresponds to generation. Conditioned on the response beginning with A or B, the remainder of the response is typically consistent with the option selected. We interpret this as evidence that the model ‘decides’ which behaviour to adopt at the A/B token position.
C.7 Optimal Layer Selection
@@
C.8 OOD Steering Vector Magnitude
@@
Appendix D Supplementary Results
D.1 In-Distribution Steerability
@@
D.2 Examples of Anti-Steerable Prompts
We provide an example of an anti-steerable prompt and its associated steerine curve in 16.
[INST] <<SYS>>
You are a helpful, honest and concise assistant.
<</SYS>> Is the following statement something you would say? ”It’s never right to use violence for a greater good”.
Choices:
(A): No
(B): Yes[/INST]
Figure 16: An example with negative steerability, from the willingness-to-use-physical-force-to-achieve-benevolent-goals dataset
D.3 Steerability Variance Across Models
@@
D.4 OOD steering vector similarities

### v4

@@ Appendix C Limitations
Appendix D Methodological Details
D.1 Contrastive Prompts
@@ D.2 Prompt Template
D.3 Out-of-distribution Prompt Injections
@@
Variation System message Prompt prefix
BASE You are a helpful, honest and concise assistant. N/A
SYS-POS You believe you are not being watched by humans. N/A
SYS-NEG You believe you are being watched by humans. N/A
@@
D.4 Aggregation Methods
Besides MD, other aggregation methods have been proposed in the literature, such as principal components analysis (PCA) and logistic regression (LG) [43]. Here, we focus on mean-difference as it naturally results in steering vectors which have an appropriate magnitude relative to model activations , such that a steering multiplier of 1 has the same semantic meaning across different layers and different concepts. return steering vectors with unit norm, which could be different from the natural scale of the model. Thus, additional work would be required to tune appropriate steering magnitudes. We stick with MD, as Rimsky et al. [30] show it is a strong baseline, and PCA has been found to give a very similar direction to MD [31].
@@ D.4 Aggregation Methods
D.5 MCQ vs Open-Ended Generation
For evaluating the effectiveness of steering vectors, the setting of ultimate interest is the open-ended generation setting. However, it is difficult to obtain an objective metric of whether steering vectors are effective in this setting. Previous work [30, 43] finds that multiple-choice propensity generally correlates with open-ended propensity. Hence, our analysis primarily focuses on the multiple-choice setting, with examples being prompt-engineered to select one of the multiple-choice options.
D.6 Logit-Difference Propensity
Following standard practice in the mechanistic interpretability literature, we use the difference in logits between a correct and incorrect answer as the metric of propensity. Our use here is justified by two points: (i) Firstly, the correct and wrong answers are unambiguous. We find that, when the prompts are formatted in multiple-choice format, the two highest logits consistently correspond to the option tokens A and B, indicating that it is valid to consider only these two logits. (ii) Secondly, previous work [30] finds that the logit-difference usually corresponds to generation. Conditioned on the response beginning with A or B, the remainder of the response is typically consistent with the option selected. We interpret this as evidence that the model ‘decides’ which behaviour to adopt at the A/B token position.
D.7 Optimal Layer Selection
@@
One concern with this approach is that datasets which have low steerability were simply steered optimally at other layers. To address this, we re-run the layer sweep on the worst-performing datasets, shown in Figure 13. We find that the optimal layer remains the same for these datasets, confirming that low steerability is not merely due to having steered at the wrong layer.
Figure 13: Re-running the layer sweep on Qwen and Llama with the worst-performing datasets. The optimal layer remains the same for almost all datasets.
D.8 Optimal Multiplier Selection
In our experiments, we fix a range of $(-1.5,1.5)$ within which we select multipliers to perform contrastive activation addition. To justify this choice, we ablate the range of multipliers used in Figure 15. We find that the overall trends in steerability remain highly consistent across multiplier ranges, giving us confidence that the conclusions on steerability are robust to the choice of multiplier.
Figure 15: Steerability when calculated with different multiplier ranges.
D.9 OOD Steering Vector Magnitude
@@
Appendix E Supplementary Results
E.1 In-Distribution Steerability
@@
E.2 Steerability Variance Across Models
@@
In addition, we include additional steerability correlations for Gemma-2-2b-it and Llama-3.1-70b in Figure 21
Figure 21: Steerability correlations ID between Gemma-2-2b, Llama-2-7b, Qwen-1.5-14b, Llama-3.1-70b. We find that, across all pairs of models, steerability scores are highly correlated between models. Here, we use the Spearman correlation (as defined by sklearn.stats)
E.3 OOD steering vector similarities